OpenStudy (hatahoooochee):

Help! I need someone to check my answer, I'll fan and medal f(x) = Square root of quantity three x plus seven. , g(x) = Square root of quantity three x minus seven. Find (f + g)(x). A. Square root of six x. B. x Square root of six. C. 3x D. Square root of quantity three x plus seven. +Square root of quantity three minus plus seven.

OpenStudy (supersmart1001):

I think this is C

OpenStudy (hatahoooochee):

I go A somehow

OpenStudy (supersmart1001):

well if u Find f(x) and g(x) so that the function can be described as y = f(g(x)). y = 4/x2+9 f(x) = x + 9 g(x) = 4/x2 f(x) = x g(x)= 4/x+9 f(x) = 1/x g(x) = 4/x+9 f(x)= 4/x2 g(x) = 9

OpenStudy (supersmart1001):

I still think C

OpenStudy (hatahoooochee):

Thank you so much